# Achronal localization and representation of the causal logic from conserved current, application to massive scalar boson

### v2

@@ 5 Covariant achronal localization out of covariant conserved current
$\pi_{\phi,\Lambda}(\Delta):=\int_{\varpi(\Delta)}\big{(}J_{0}(\phi;\tau(x),x)-J(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x\>.$ (2)
As the integrand is nonnegative, $\pi_{\phi,\Lambda}$ is a $\sigma$-additive measure. By (a) and (4.3), $\pi_{\phi,\Lambda}(\Lambda)=||\phi||^{2}$. Due to (5), (5) $\phi\mapsto\pi_{\phi,\Lambda}(\Delta)$ is the quadratic form of a bounded Hermitian sesquilinear form on $\mathcal{D}$. Therefore by [12, Lemma 48] there is a bounded operator $T(\Delta)$, $0\leq T(\Delta)\leq I$ with $\pi_{\phi,\Lambda}(\Delta)=\langle\phi,T_{\Lambda}(\Delta)\phi\rangle$.
       Let $\phi\in\mathcal{H}$. By continuity on $\phi$, $\Delta\mapsto\langle\phi,T_{\Lambda}(\Delta)\phi\rangle$ is finitely additive. Actually it is $\sigma$-additive. Indeed, let $\Delta_{n}\downarrow_{n}\emptyset$ for Borel $\Delta_{n}\subset\Lambda$. Let $\epsilon>0$. Let $\phi^{\prime}\in\mathcal{D}$ with $||\phi-\phi^{\prime}||\leq\epsilon$. An obvious application of the triangle inequality yields $||\langle\phi,T_{\Lambda}(\Delta_{n})\phi\rangle-\langle\phi^{\prime},T_{\Lambda}(\Delta_{n})\phi^{\prime}\rangle||\leq||\phi-\phi^{\prime}||\,||\phi||+||\phi^{\prime}||\,||\phi-\phi^{\prime}||\leq 2\epsilon||\phi||+\epsilon^{2}\leq C\epsilon$ with $C<\infty$ independent of $\Delta_{n}$. As $\langle\phi^{\prime},T_{\Lambda}(\Delta_{n})\phi^{\prime}\rangle\to 0$, one infers $\langle\phi,T_{\Lambda}(\Delta_{n})\phi\rangle\to 0$, whence the claim.
       So $T_{\Lambda}$ is weakly $\sigma$-additive, which by [21, Theorem 4.28] implies the strong $\sigma$-additivity.
@@
       (a) the achronal set $g\cdot\Delta$ equals $\{(\tau_{g}(y),y):y\in\varpi(g\cdot\ \Delta)\}$ for $\tau_{g}(y):=(g\cdot(\tau(x),x))_{0}$ with $x:=S^{-1}(y)$, where $S:\mathbb{R}^{3}\to\mathbb{R}^{3}$, $S(x):=\varpi\big{(}g\cdot(\tau(x),x)\big{)}$ is a bijection.
@@
       (b) By (5) (ii), $\pi_{W(g)^{-1}\phi,\Lambda}(\Delta)=\int_{\varpi(\Delta)}\mathfrak{J}\big{(}\phi,g\cdot(\tau(x),x)\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(x))\big{)}\operatorname{d}^{3}x=\int_{\varpi(\Delta)}\mathfrak{J}\big{(}\phi,\tau_{g}(S(x)),S(x)\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(x))\big{)}\operatorname{d}^{3}x$, which using the image of the Lebesgue measure $\lambda$ equals $\int_{S(\varpi(\Delta))}\mathfrak{J}\big{(}\phi,\tau_{g}(y),y\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(S^{-1}(y))\big{)}\operatorname{d}S(\lambda)(y)$. Now recall $S(\varpi(\Delta))=\varpi(g\cdot\Delta)$ and note $\operatorname{d}S(\lambda)/\operatorname{d}\lambda=|\det D\,S^{-1}|=|\det D\,S(S^{-1}(\cdot))|^{-1}$.
@@
which is easy in the case $A\in SU(2)$. So it suffices to check the case $g=\operatorname{e}^{\rho\sigma_{3}/2}$, $\rho\in\mathbb{R}$444 Explicitly $\operatorname{e}^{\,\rho\,\sigma_{3}/2}=\operatorname{diag}(\operatorname{e}^{\rho/2},\operatorname{e}^{-\rho/2})$ acts on $\mathbb{R}^{4}$ by $\left(\begin{array}[]{cccc}\cosh(\rho)&0&0&\sinh(\rho)\\ 0&1&0&0\\ 0&0&1&0\\ \sinh(\rho)&0&0&\cosh(\rho)\end{array}\right)$ for $\rho\in\mathbb{R}$ Put $c:=\cosh\rho$, $s:=\sinh\rho$, $z:=\nabla\tau(x)$, $x=S^{-1}(y)$. The rows of $\big{(}D\,S(x)\big{)}^{-1}$ are $(1,0,0)$, $(0,1,0)$, $\frac{1}{c+sz_{3}}(-sz_{1},-sz_{2},1)$. So the right side of (*) equals $\frac{1}{c+sz_{3}}(c+sz_{3},z_{1},z_{2},cz_{3}+s)$. On the left hand side $\nabla\tau_{g}(y))=\big{(}cz_{1}-\frac{(cz_{3}+s)sz_{1}}{c+sz_{3}},\dots,\frac{cz_{3}+s}{c+sz_{3}}\big{)}$. Hence (*) holds thus accomplishing the proof.∎
@@ 6 Covariant achronal localizations of the massive scalar boson
       To be explicit in the following we use the momentum representation with $L^{2}(\mathbb{R}^{3})$ being the space of states. By [6, (6.1), (11) Theorem] one has
@@
       The aim is to extend $T$ to an AL which is Poincaré covariant under the representation $W$ describing the massive scalar boson [^5] Often one uses the antiunitarily equivalent $\operatorname{e}^{-\operatorname{i}\mathfrak{a}\cdot\,\mathfrak{p}}$.
  • $\big{(}W(\mathfrak{a},A)\phi\big{)}(p)=\sqrt{\epsilon(q)/\epsilon(p)}\,\operatorname{e}^{\operatorname{i}\mathfrak{a}\cdot\,\mathfrak{p}}\,\phi(q)$
    with $\epsilon(p):=\sqrt{m^{2}+p^{2}}$, $\mathfrak{p}:=(\epsilon(p),p)$, $\mathfrak{q}=(q_{0},q):=A^{-1}\cdot\mathfrak{p}$
following the considerations in sec. 5. Petzold and collaborators [14] show that $J_{0}$ is the zero component of a covariant conserved four-vector current $\mathfrak{J}:=(J_{0},J)$ if and only if
@@
where $g:[m^{2},\infty[\to\mathbb{R}$ is continuous with $g(m^{2})=1$ such that the zeroth component $K_{0}$ of $\mathfrak{K}$ is a positive definite kernel on $\mathbb{R}^{3}$ (see also [6, (55) Corollary]).
@@
Proof. Let $\alpha$ be a multiindex with $|\alpha|\leq 4$. Note $x^{\alpha}\operatorname{e}^{\operatorname{i}px}=(-\operatorname{i})^{|\alpha|}D_{p}^{\alpha}\operatorname{e}^{\operatorname{i}px}$. Multiply $J_{0}(\phi;0,x)$ in (6.1) by $x^{\alpha}$. After integration by parts the integrand is bounded by $\big{|}D_{p}^{\alpha}\big{(}K_{0}(k,p)\phi(p)\big{)}\big{|}\,|\phi(k)|$, which is integrable. So $|x^{\alpha}J_{0}(\phi;0,x)|$ is bounded by a constant independent of $x$.∎
@@
(24) Theorem. Let $\mathfrak{J}$ be a covariant conserved current with causal kernel for the massive scalar boson. Assume $g\in C^{4}([m^{2},\infty[)$. Then there is an AL $T$ satisfying for every achronal Borel set $\Delta$ and $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$
@@
We turn to the family of localizations of the massive scalar boson obtained out of its stress energy tensor [18]. For a thorough treatment see [12, sec. 6]. The family is indexed by the normalized future-directed timelike fourvectors $\mathfrak{n}$, i.e. $\mathfrak{n}^{\cdot 2}=1$, $n_{0}>0$. The related currents read still in the momentum representation with $\phi\in C_{c}^{\infty}(\mathbb{R}^{3})$
@@
       One easily checks that $\mathfrak{J}_{\mathfrak{n}}(\phi,\cdot)$ is smooth bounded conserved and that $J_{\mathfrak{n},0}(\cdot\,;0,x)$ is quadratic for every $x\in\mathbb{R}^{3}$. Also one verifies the covariance
@@
for all $\mathfrak{n}$, $g=(\mathfrak{a},A)$, $\mathfrak{x}$. In addition, for every $\mathfrak{n}$, $\phi$, one has $J_{\mathfrak{n},0}(\phi,;0,\cdot)\geq 0$ and $\int J_{\mathfrak{n},0}(\phi,;0,x)\operatorname{d}^{3}x=||\phi||^{2}$ as shown in [12, (64), Theorem 54]. Lemma (5), (5) hold for every $\mathfrak{J}_{\mathfrak{n}}$ by the same proofs due to (6.5). Finally, (5)(b) holds for every $J_{\mathfrak{n},0}$ by a proof analogous to that of (6). In summary, (5) applies to $\mathfrak{J}_{\mathfrak{n}}$. It follows
@@ 7 Covariant representation of the causal logic for the massive scalar boson
       The causal logic is thoroughly studied by Cegla, Jadczyk, Jancewicz in [9] and has been studied further on. $\mathcal{C}$ enjoys some appealing physical properties similar to the ones of the lattice of abstract elementary propositions of a quantum system [9] and see also the more recent works [3, 4]: it is possible to prove that the lattice $\mathcal{C}$ is $\sigma$-complete, irreducible, orthomodular, atomic, atomistic exactly as for a quantum lattice, but fails to satisfy the covering law and separability.
@@
Only recently covariant RCL have been constructed. They concern quantum mechanical systems with definite spin $j\in\mathbb{N}_{0}/2$ and mass spectrum $\subset$ ]0,[]0$\infty$[] 0 , ∞ [ of positive Lebesgue measure [7]. Moreover, due to the results in [5], there is evidence that there are also RCL regarding the Dirac particles and the four Weyl fermions. These results have been obtainded by group theoretical methods.
@@
       Recently an other idea of localization of the massive scalar boson in causally complete regions is pursued [19]. It uses the method of modular localization from AQFT. Given a Cauchy surface, via the modular localization map, to every state and to the causal completion of every its Borel subsets a probability of localization is attributed. This is asymptotically additive when increasing the distance between the regions of localization.
@@
(28) Theorem. To every causal kernel $\mathfrak{K}$ with $g\in C^{4}([m^{2},\infty[)$ there is a unique covariant RCL $F$ for the massive scalar boson such that

### v4

@@ 5 Covariant achronal localization out of covariant conserved current
$\pi_{\phi,\Lambda}(\Delta):=\int_{\varpi(\Delta)}\big{(}J_{0}(\phi;\tau(x),x)-J(\phi;\tau(x),x)\nabla\tau(x)\big{)}\operatorname{d}^{3}x$ (2)
As the integrand is nonnegative, $\pi_{\phi,\Lambda}$ is a $\sigma$-additive measure. By (a) and (5), $\pi_{\phi,\Lambda}(\Lambda)=||\phi||^{2}$. Due to (5), (5), $\phi\mapsto\pi_{\phi,\Lambda}(\Delta)$ is the quadratic form of a bounded Hermitian sesquilinear form on $\mathcal{D}$. Therefore by [13, Lemma 48] there is a bounded operator $T(\Delta)$, $0\leq T(\Delta)\leq I$ with $\pi_{\phi,\Lambda}(\Delta)=\langle\phi,T_{\Lambda}(\Delta)\phi\rangle$.
       Let $\phi\in\mathcal{H}$. Approximating $\phi$ by $\phi_{n}\in\mathcal{D}$ finite additivity of $\Delta\mapsto\langle\phi,T_{\Lambda}(\Delta)\phi\rangle$ follows by continuity. Actually it is $\sigma$-additive. Indeed, let $\Delta_{n}\downarrow_{n}\emptyset$ for Borel $\Delta_{n}\subset\Lambda$. Let $\epsilon>0$. Let $\phi^{\prime}\in\mathcal{D}$ with $||\phi-\phi^{\prime}||\leq\epsilon$. An obvious application of the triangle inequality yields $||\langle\phi,T_{\Lambda}(\Delta_{n})\phi\rangle-\langle\phi^{\prime},T_{\Lambda}(\Delta_{n})\phi^{\prime}\rangle||\leq||\phi-\phi^{\prime}||\,||\phi||+||\phi^{\prime}||\,||\phi-\phi^{\prime}||\leq 2\epsilon||\phi||+\epsilon^{2}\leq C\epsilon$ with $C<\infty$ independent of $\Delta_{n}$. As $\langle\phi^{\prime},T_{\Lambda}(\Delta_{n})\phi^{\prime}\rangle\to 0$, one infers $\langle\phi,T_{\Lambda}(\Delta_{n})\phi\rangle\to 0$, whence the claim.
       So $T_{\Lambda}$ is weakly $\sigma$-additive, which by [27, Theorem 4.28] implies the strong $\sigma$-additivity.
@@
       (a) the achronal set $g\cdot\Delta$ is given by
$g\cdot\Delta=\{(\tau_{g}(y),y):y\in\varpi(g\cdot\ \Delta)\}$
for $\tau_{g}(y):=(g\cdot(\tau(x),x))_{0}$ with $x:=S^{-1}(y)$, where
$S:\mathbb{R}^{3}\to\mathbb{R}^{3},\quad S(x):=\varpi\big{(}g\cdot(\tau(x),x)\big{)}$
is a bijection.
@@
       (b) By (5) (ii),
$\pi_{W(g)^{-1}\phi,\Lambda}(\Delta)=\int_{\varpi(\Delta)}\mathfrak{J}\big{(}\phi,g\cdot(\tau(x),x)\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(x))\big{)}\operatorname{d}^{3}x$
equals $\int_{\varpi(\Delta)}\mathfrak{J}\big{(}\phi,\tau_{g}(S(x)),S(x)\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(x))\big{)}\operatorname{d}^{3}x$ which using the image of the Lebesgue measure $\lambda$ becomes
$\int_{S(\varpi(\Delta))}\mathfrak{J}\big{(}\phi,\tau_{g}(y),y\big{)}\cdot\big{(}A\cdot(1,\nabla\tau(S^{-1}(y))\big{)}\operatorname{d}S(\lambda)(y)$
Now recall $S(\varpi(\Delta))=\varpi(g\cdot\Delta)$ and note $\operatorname{d}S(\lambda)/\operatorname{d}\lambda=|\det D\,S^{-1}|=|\det D\,S(S^{-1}(\cdot))|^{-1}$.
@@
which is easy in the case $A\in SU(2)$. So it suffices to check the case $g=\operatorname{e}^{\rho\sigma_{3}/2}$, $\rho\in\mathbb{R}$555 Explicitly $\operatorname{e}^{\,\rho\,\sigma_{3}/2}=\operatorname{diag}(\operatorname{e}^{\rho/2},\operatorname{e}^{-\rho/2})$ acts on $\mathbb{R}^{4}$ by $\left(\begin{array}[]{cccc}\cosh(\rho)&0&0&\sinh(\rho)\\ 0&1&0&0\\ 0&0&1&0\\ \sinh(\rho)&0&0&\cosh(\rho)\end{array}\right)$ for $\rho\in\mathbb{R}$ Put $c:=\cosh\rho$, $s:=\sinh\rho$, $z:=\nabla\tau(x)$, $x=S^{-1}(y)$. The rows of $\big{(}D\,S(x)\big{)}^{-1}$ are $(1,0,0)$, $(0,1,0)$, $\frac{1}{c+sz_{3}}(-sz_{1},-sz_{2},1)$. So the right side of (*) equals $\frac{1}{c+sz_{3}}(c+sz_{3},z_{1},z_{2},cz_{3}+s)$. On the left hand side $\nabla\tau_{g}(y))=\big{(}cz_{1}-\frac{(cz_{3}+s)sz_{1}}{c+sz_{3}},\dots,\frac{cz_{3}+s}{c+sz_{3}}\big{)}$. Hence (*) holds thus accomplishing the proof.∎
@@ 6 Covariant achronal localizations of the massive scalar boson
The theory developed so far, in particular the construction of a covariant AL (5), (5), is rather general without requiring additional effort. It refers to any relativistic quantum mechanical system, which is completely determined by a unitary representation $W$ of the Poincaré group. Here we turn to the elementary particle being the massive scalar boson, which is the one-particle system of a free Klein-Gordon quantum field uniquely determined by $W=[m,0,+]$, i.e., the mass $m>0$, the spin 00 and the sign of energy $+$. In the momentum representation with $L^{2}(\mathbb{R}^{3})$ being the space of states $W$ reads [^6] Often one uses the antiunitarily equivalent $\operatorname{e}^{-\operatorname{i}\mathfrak{a}\cdot\,\mathfrak{p}}$.
$\big{(}W(\mathfrak{a},A)\phi\big{)}(p)=\sqrt{\epsilon(q)/\epsilon(p)}\,\operatorname{e}^{\operatorname{i}\mathfrak{a}\cdot\,\mathfrak{p}}\,\phi(q)$ (6.1)
with $\epsilon(p):=\sqrt{m^{2}+p^{2}}$, $\mathfrak{p}:=(\epsilon(p),p)$, $\mathfrak{q}=(q_{0},q):=A^{-1}\cdot\mathfrak{p}$.
@@
       In the momentum representation by [7, (6.1), (11) Theorem] one has
@@
       The aim is to extend $T$ to an AL which is Poincaré covariant under the representation $W$ (6.1) of the massive scalar boson following the considerations in sec. 5. Petzold and collaborators [15] show that $J_{0}$ is the zero component of a covariant conserved four-vector current $\mathfrak{J}:=(J_{0},J)$ if and only if
@@
where $g:[m^{2},\infty[\to\mathbb{R}$ is continuous with $g(m^{2})=1$ such that the zeroth component $K_{0}$ of $\mathfrak{K}$ is a positive definite kernel on $\mathbb{R}^{3}$ (see also [7, (55) Corollary]).
@@
Proof. Let $x_{0}\neq 0$. Put
$F:\mathbb{R}^{3}\times\mathbb{R}^{3}\to\mathbb{R},\quad F(k,p):=\frac{\varepsilon(k)x_{0}-kx}{|x|+|x_{0}|}-\frac{\varepsilon(p)x_{0}-px}{|x|+|x_{0}|}$
$F$ is $C^{\infty}$. Put $\varphi(k,p):=(2\pi)^{-3}K_{0}(k,p)\overline{\phi(k)}\phi(p)$. $\varphi$ is $C_{c}^{4}$. One has
$J_{0}(\phi,\mathfrak{x})=\int\int\operatorname{e}^{\operatorname{i}(|x|+|x_{0}|)F(k,p)}\varphi(k,p)\operatorname{d}^{3}k\operatorname{d}^{3}p$ (*)
We proceed according the proof of [26, Theorem 1.8], which concerns the non-stationary phase method. A positive lower bound (**) of $|\nabla F|$ is crucial.
       Let $K:=\operatorname{supp}\phi$. Put $\beta:=\max\{\frac{|p|}{\varepsilon(p)}:p\in K\}$. Clearly $10\leq\beta<$ 1. Then
$\nabla_{k}F(k,p)=(|x|+|x_{0}|)^{-1}\big{(}\frac{x_{0}}{\epsilon(k)}k-x\big{)}$
whence $|\nabla_{k}F(k,p)|\geq(|x|+|x_{0}|)^{-1}\big{(}|x|-|x_{0}|\frac{|k|}{\epsilon(k)}\big{)}\geq\frac{|x|-\beta|x_{0}|}{|x|+|x_{0}|}$ for $k\in K$. Now assume $|x|\geq|x_{0}|$. Then $|\nabla_{k}F(k,p)|\geq\frac{1-\beta}{2}$ and similarly $|\nabla_{p}F(k,p)|\geq\frac{1-\beta}{2}$ for $p\in K$. It follows
$|\nabla F(k,p)|\geq\frac{1-\beta}{\sqrt{2}}>0\;\text{ for }|x|\geq|x_{0}|,\;(k,p)\in K\times K$ (**)
Note also that the derivatives satisfy $|D_{k}^{\alpha}F(k,p)|\leq 1$, $|D_{p}^{\alpha}F(k,p)|\leq 1$ for $|\alpha|=1$, $|D_{k}^{\alpha}F(k,p)|\leq|D^{\alpha}\varepsilon(k)|$, $|D_{p}^{\alpha}F(k,p)|\leq|D^{\alpha}\varepsilon(p)|$ for $|\alpha|\geq 2$. Moreover, $\operatorname{supp}\varphi\subset K\times K$.
       Repeated integration by parts as in the proof of [26, Theorem 1.8] yields
$J_{0}(\phi,\mathfrak{x})=(|x|+|x_{0}|)^{-n}\int_{K^{2}}\operatorname{e}^{\operatorname{i}(|x|+|x_{0}|)F(v)}\Phi^{(n)}(v)\operatorname{d}^{6}v\;\text{ for }|x|\geq|x_{0}|,\,n\leq 4$
using the notation $v:=(k,p)\in\mathbb{R}^{6}$ and $K^{2}=K\times K\subset\mathbb{R}^{6}$. Here the function $\Phi^{(n)}$ is a sum of products of factors $|\nabla F|^{-2}$, $D^{\alpha}F$ for $1\leq|\alpha|\leq 4$, $D^{\alpha}\varphi$ for $|\alpha|\leq 4$. Hence $\Phi^{(n)}$ is bounded on $K^{2}$ independent of $|x|,|x_{0}|$. The result follows.∎
@@
(24) Theorem. Let $\mathfrak{J}$ be a covariant conserved current 777recall (5) with causal kernel for the massive scalar boson. Assume $g\in C^{4}([m^{2},\infty[)$. Then there is an AL $T$ satisfying for every achronal Borel set $\Delta$ and $\phi\in C^{\infty}_{c}(\mathbb{R}^{3})$
@@
We turn to the family of localizations of the massive scalar boson obtained out of its stress energy tensor [23]. For a thorough treatment see [13, sec. 6]. The family is indexed by the normalized future-directed timelike fourvectors $\mathfrak{n}$, i.e. $\mathfrak{n}^{\cdot 2}=1$, $n_{0}>0$. The related currents read still in the momentum representation with $\phi\in C_{c}^{\infty}(\mathbb{R}^{3})$
@@
       One easily checks that $\mathfrak{J}_{\mathfrak{n}}(\phi,\cdot)$ is real smooth bounded conserved and that $J_{\mathfrak{n},0}(\cdot\,;0,x)$ is quadratic for every $x\in\mathbb{R}^{3}$. Also one verifies the covariance
@@
for all $\mathfrak{n}$, $g=(\mathfrak{a},A)$, $\mathfrak{x}$. In addition, for every $\mathfrak{n}$, $\phi$, one has $J_{\mathfrak{n},0}(\phi,;0,\cdot)\geq 0$ and $\int J_{\mathfrak{n},0}(\phi,;0,x)\operatorname{d}^{3}x=||\phi||^{2}$ as shown in [13, (64), Theorem 54]. Lemma (5), (5), and (5) hold for every $\mathfrak{J}_{\mathfrak{n}}$ by the same proofs due to (6.6). Finally, (5)(b) holds for every $J_{\mathfrak{n},0}$ by a proof analogous to that of (6). In summary, (5) applies to $\mathfrak{J}_{\mathfrak{n}}$. It follows
@@ 7 Covariant representation of the causal logic for the massive scalar boson
       The causal logic is thoroughly studied by Cegla, Jadczyk, in [10] and has been studied further on. $\mathcal{C}$ enjoys some appealing physical properties similar to the ones of the lattice of abstract elementary propositions of a quantum system [10] and see also the more recent works [4, 5]: it is possible to prove that the lattice $\mathcal{C}$ is $\sigma$-complete, irreducible, orthomodular, atomic, atomistic exactly as for a quantum lattice, but fails to satisfy the covering law and separability.
@@
Only recently covariant RCL have been constructed. They concern quantum mechanical systems with definite spin $j\in\mathbb{N}_{0}/2$ and mass spectrum $\subset$ ]0,[]0$\infty$[] 0 , ∞ [ of positive Lebesgue measure [8]. Moreover, due to the results in [6], there is evidence that there are also RCL regarding the Dirac particles and the four Weyl fermions. These results have been obtained by group theoretical methods.
@@
       Recently an other idea of localization of the massive scalar boson in causally complete regions is pursued [24]. It uses the method of modular localization from AQFT. Given a Cauchy surface, via the modular localization map, to every state and to the causal completion of every of its Borel subsets a probability of localization is attributed. This is asymptotically additive when increasing the distance between the regions of localization.
@@
(28) Theorem. To every causal kernel $\mathfrak{K}$ with $g\in C^{4}([m^{2},\infty[)$ there is a unique covariant888recall (7) RCL $F$ for the massive scalar boson such that
@@
8 Final comments and QFT outlook
All achievements presented in this work concern the notion of localization for a boson massive particle in Minkowski spacetime. We constructed very general notions of achronal localization in complete agreement with requirements of Poincaré covariance and causality. It was thanks to suitable probability conserved currents and an appropriate use of the divergence theorem. Referring to the discussion in sec.5, the first type of causal obstructions against any localization notion, historically arising from Hegerfeldt’s analysis and considered here into a more modern perspective started in [6], were found to be definitely harmless. A fully-fledged notion of achronal, causal, covariant localization is proved to be possible, provided one adopts the modern point of view where, in the general case, observables are described by (families of) POVMs instead of (families of) projection-valued measures, namely selfadjoint operators. No sharply localized probability distributions are permitted in this approach. (Though sharply distributions can be approximated with arbitrary precision [6, 23].) It is interesting noting that, classic notions as the Newton-Wigner position operator, apparently ruled out by the Hegerfeldt theorem, still keep to play a role in our framework as the first-moment of the introduced families of POVMs as established in [13] in a wide generality. This fact gives rise to the standard notion of localization in the rest space of an observer under suitable non-relativistic conditions (large mass limit).
Two types of probability currents, associated to respective notions of localization, have been considered in this work for real massive bosons (see sec.6). A type, extensively studied in [7] within a general framework, was based on the theory of causal kernels. Another type discussed in [23, 13], was constructed in terms of the stress-energy tensor.
The new perspective developed in this paper on the one hand refers to a completely general notion of rest space [7] where localization takes place. This notion is even more general than the one studied in [13], and is represented by (maximal) achronal sets, using the weakest possible regularity condition explicitly required by physics. On the other hand, this generality permitted us to construct, apparently for the first time, an effect-space covariant representation of the lattice of causally complete Borel sets of the spacetime, the logic of spacetime, for a boson particle with definite and strictly-positive value of the mass. A similar type of representation for particles with unsharp mass spectrum was already introduced in [7]. In this view, a particle localized in an (roughly speaking $3D$) achronal region $\Delta$ is the same as a particle localized in its (roughly speaking $4D$) causal completion $\Delta^{\sim}$. Or also in every other achronal region $\Delta^{\prime}$ whose causal completion is the same $\Delta^{\sim}$.
From the perspective of QFT, our approach concerns the one-particle structure of the Fock representation built upon the Poincaré invariant Gaussian vacuum state in Minkowski spacetime, for a massive real Klein-Gordon field. As a consequence, at this stage of the study, it is not directly possible to implement a fully-fledged analysis of local causality in the Haag-Kastler perspective. An evident issue concerns the fact that the projector onto the one-particles Hilbert space does not belong to a local algebra of observables. To overcome this obstruction it is necessary to see all introduced notions as restrictions to the one-particle space of more general notions defined in terms of local algebras of the Klein-Gordon field. The restrictions should be considered as imposed by the choice of the used one-particle states.
Referring to the two types of localizations analyzed in sec.6, this is surely possible concerning the notion of localization constructed out the stress energy tensor [13]. This is because the used stress-energy tensor operator is nothing but the one of the full quantum field averaged against one-particle states. The notion arising from causal kernels [7] needs a deeper analysis. However, it seems clear that, at least in the case of a charged field, a relation with the standard conserved current of the whole quantum field must exist.
Yet, the QFT perspective is the right place where tackling the second type of causal obstructions (see sec.5) against any notion of localization in addition to Hegerfeldt’s theorems. We are referring to the already mentioned Malament [22] and Halvorson-Clifton [16] theorems. These no-go results show that localization operators (either orthogonal projectors or effects) associated to causally separated regions $\Delta$ and $\Delta^{\prime}$ cannot commute, under physically plausible assumptions, as additivity and a bound from below of the generator of time translations. The commutativity requirement as mathematical description of causal independence is an indispensable condition in the local algebras formulation.
All these issues will be tackled elsewhere.